Machinery purchased for $61,800 by Swifty Co. in 2016 was originally estimated to have a life of 8 years with a salvage value of $4,120 at the end of that time. Depreciation has been entered for 5 years on this basis. In 2021, it is determined that the total estimated life should be 10 years with a salvage value of $4,635 at the end of that time. Assume straight-line depreciation.

Prepare the entry to correct the prior years' depreciation, if necessary. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.) Account Titles and Explanation Debit Credit enter an account title enter a debit amount enter a credit amount enter an account title enter a debit amount enter a credit amount

Prepare the entry to record depreciation for 2021. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Use Machinery related account.) Account Titles and Explanation Debit Credit enter an account title enter a debit amount enter a credit amount enter an account title enter a debit amount enter a credit amount
Answer #1

a) Journal entry

Date account and explanation Debit Credit
No Journal entry

b) Accumulated depreciation = (61800-4120)/8*5 = 36050

2021 Depreciation = (61800-36050-4635)/5 = 4223

Date account and explanation Debit Credit
Depreciation expense 4223
Accumulated depreciation-equipment 4223